The length of a rectangular garden is 3 meters greater than the width. The area of the gard4en is 108 square meters. Find the length.
State your answer in terms of meters, but do not include a unit of measure with your response. Although you may find the width, do not include this information in the answer?
: The length of a rectangular garden is 3 meters greater than the width. The area of the gard4en is 108 square meters. Find the length.
State your answer in terms of meters, but do not include a unit of measure with your response. Although you may find the width, do not include this information in the answer?

Answer by vleith(1174)

Given: Length = Width + 3
Width = Length - 3
Area = Length*Width
Area = 108
A = L*W
180 = L * (L-3)
180 = L^2 - 3L
0 = L^2 - 3L - 180
0 = (L -15)(L + 12)
Either or both terms must be 0 for the product to equal 0.
Either (L-15) = 0 or (L+12)=0
So L = 15 or L = -12.
Since a length cannot be negative, the length must be 15.